A project has an NPV of $44,000. Calculate the cost of capital of this project if it generates the following cash flows for six years after an initial investment of $190,000: (Round answer to 2 decimal places, e.g. 25.25%.)

Year 1: $44,000
Year 2: $44,000
Year 3: $25,000
Year 4: $72,000
Year 5: $56,000
Year 6: $67,000

cost of capital= ?%
Answer #1

NPV=Present value of inflows-Present value of outflows

Present value of inflows=(44000+190,000)=$234000

Let cost of capital be x%

At this rate;Present value of inflows=cash inflow*Present value of discounting factor(rate%,time period)

234000=44000/1.0x+44000/1.0x^2+25000/1.0x^3+72000/1.0x^4+56000/1.0x^5+67000/1.0x^6

Hence x=cost of capital=7.69%(Approx)